A 1.8 kg otter starts from rest at the top of a
muddy incline 83.5 cm long and slides down
to the bottom in 0.50 s.
What net external force acts on the otter
along the incline?
Answer in units of N.
1
Expert's answer
2019-12-23T11:29:17-0500


By the condition of the problem "s=AB=83.5sm=0.835m"

from the formula "s=\frac{at^2}{2}" we determine the acceleration

"a=\frac{2s}{t^2}=\frac{2\cdot 0.835}{0.5^2}=6.68{m/s^2}"

then we write down Newton’s second law

"F=m \cdot a=1.8 \cdot 6.68=12.024N"